# BiMediX2 : Bio-Medical EXpert LMM for Diverse Medical Modalities

### v1

@@ Abstract
This paper introduces BiMediX2, a bilingual (Arabic-English) Bio-Medical EXpert Large Multimodal Model (LMM) with a unified architecture that integrates text and visual modalities, enabling advanced image understanding and medical applications. BiMediX2 leverages the Llama3.1 architecture and integrates text and visual capabilities to facilitate seamless interactions in both English and Arabic, supporting text-based inputs and multi-turn conversations involving medical images. The model is trained on an extensive bilingual healthcare dataset consisting of 1.6M samples of diverse medical interactions for both text and image modalities, mixed in Arabic and English. We also propose the first bilingual GPT-4o based medical LMM benchmark named BiMed-MBench. BiMediX2 is benchmarked on both text-based and image-based tasks, achieving state-of-the-art performance across several medical benchmarks. It outperforms recent state-of-the-art models in medical LLM evaluation benchmarks. Our model also sets a new benchmark in multimodal medical evaluations with over 9% improvement in English and over 20% in Arabic evaluations. Additionally, it surpasses GPT-4 by around 9% in UPHILL factual accuracy evaluations and excels in various medical Visual Question Answering, Report Generation, and Report Summarization tasks. The project page, including source code and the trained model, is available at  https://github.com/mbzuai-oryx/BiMediX2.
@@ 1 Introduction
Figure 1: Model Performance Comparison on BiMed-MBench: These comparisons are made across different categories, including CT, MRI, CXR, Histology, Gross, and their Arabic counterparts (CT_ar, MRI_ar, CXR_ar, Histology_ar, Gross_ar). The models compared are LLaVA-$p$, LLaVA-Med, BiMediX2, Dragonfly-Med, MiniGPT-Med, and BiomedGPT. Each axis represents the performance score in a specific category, allowing for a visual comparison of how each model performs in bilingual medical contexts.
Recent advancements in healthcare AI, including medical large language models (LLMs) and large multimodal models (LMMs), have shown great potential in improving access to quality medical advice. However, these advancements remain predominantly English-centric, leaving significant gaps in addressing the healthcare needs of non-English-speaking populations. This language bias limits the accessibility and utility of AI-driven medical solutions in regions where languages like Arabic are widely spoken, creating a pressing need for more inclusive and multilingual models. Furthermore, conventional medical LMMs (Zhang et al. (2024); Li et al. (2023); Chen et al. (2024) see Fig 3) generally compromise their advanced medical text-based understanding when integrating multimodal capabilities.
Our project, BiMediX2, strives to address the aforementioned limitations of existing medical LMMs by introducing a Bio-Medical EXpert Large Multimodal Model (LMM) built on the Llama3.1 architecture. It achieves state-of-the-art performance in medical image understanding while also excelling in text-based medical LLM evaluations. BiMediX2 facilitates seamless user interactions in both English and Arabic, leveraging our extensive, purpose-built bilingual multimodal healthcare dataset BiMed-V.
BiMediX2 facilitates medical image understanding along with textual-queries, enabling multi-turn conversations about user-provided medical images. It supports a wide range of imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images. This multimodal approach ensures comprehensive and accessible medical assistance in a bilingual setting.
Recent advancements in medical LLMs and vision-language models (VLMs) have shown progress in tasks such as multi-turn conversations (MTC), report generation (RG), and specialized domains like radiology (Rad) and pathology (Path) as shown in Tab 1. While models like LLaVa-pp (Rasheed et al. (2024)) and LLaVa-Med (Li et al. (2023)) enable seamless multi-turn conversations, others, such as MiniGPT-Med (Alkhaldi et al. (2024)), are limited to specific modalities like chest X-rays, and BiomedGPT (Zhang et al. (2024)) requires fine-tuned checkpoints for each downstream task. BiMediXv1 (Pieri et al. (2024)) enables multi-turn conversational capabilities and report summarization within a bilingual framework, setting it apart from other medical LLMs. A comprehensive comparison of state-of-the-art models in Clinical LLM evaluations is shown in Fig 3.
Building on this, BiMediX2 expands its functionality to cover a broader spectrum of medical tasks, including Multi-turn Conversations (MTC), Report Summarization (RS), Report Generation (RG), Radiology (Rad), Ophthalmology (Oph), Pathology (Path), Microscopic Analysis (Micro), in a Unified Model (UM) with LLM and VLM capabilities (LLM+VLM). Fig 1 illustrates the performance of state-of-the-art medical LMMs on our BiMed-MBench evaluation in a bilingual context. Its robust bilingual support for Arabic addresses the healthcare needs of Arabic-speaking regions, providing a comprehensive and inclusive solution for diverse, multilingual, and multimodal healthcare applications.
@@
Table 1: Comparison of Recent Medical LLMs and VLMs. Abbreviations: MTC (Multi-turn conversation), RS (Report Summarization), RG (Report Generation), Rad (Radiology), Oph (Ophthalmology), Path (Pathology), Micro (Microscopic), UM (Unified Model: Single model checkpoint for all downstream tasks), LLM+VLM (Unified LLM + VLM), Bil (Ar) (Bilingual Arabic capabilities).
Contributions
In this work, we present BiMediX2, a bilingual multimodal medical LMM with the following key contributions:
(i) We introduce the first bilingual medical LMM that achieves state-of-the-art results on VLM evaluation benchmarks across various medical image modalities, while also excelling on medical LLM evaluation benchmarks.
(ii) We curated a comprehensive Arabic-English multimodal bilingual instruction set named BiMed-V comprising over 1.6M instructions.
@@
(iii) We introduce the first bilingual GPT-4o-based medical LMM benchmark named BiMed-MBench, consisting of 286 medical queries in English and Arabic across various medical image modalities, fully verified by medical experts.
(iv) Our BiMediX2 LLM outperforms GPT-4 by more than 8% on the USMLE benchmark and by more than 9% in UPHILL factual accuracy evaluations.
@@
(v) Our BiMediX2 LMM achieves state-of-the-art results on BiMed-MBench, with over a 9% improvement in English evaluations and more than a 20% improvement in Arabic evaluations. Furthermore, it excels in medical Visual Question Answering, Report Generation, and Report Summarization tasks.
@@ 2 BiMediX2
Figure 2: BiMediX2: Overall Architecture Our model is designed for medical image analysis and bilingual multi-turn conversations. Medical images are processed through a Vision Encoder and aligned with a Projector, while the text inputs are tokenized using the default tokenizer. The resulting tokens are then passed into the language model (Meta Llama 3.1) to generate responses in the prompted language. We only train the language model using LoRA adapters, while the projector is finetuned for medical image-text alignment. A robust data generation framework translates an English data corpus into Arabic using GPT-4o, with verification by a medical expert to ensure accurate and contextually appropriate translations. This approach supports effective training and benchmarking in a bilingual context.
The architecture of BiMediX2 is meticulously designed to facilitate seamless integration of medical image analysis and bilingual multi-turn conversations. At its core, (see Fig 2) the model employs a Vision Encoder (Radford et al. (2021)) to process a diverse array of medical imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images. This visual data is aligned with textual inputs through a dedicated Projector, ensuring accurate and contextually rich medical image-text mapping. Text inputs are processed using a standard tokenizer, transforming them into the language embedding space of Llama 3.1 (Meta (2024)). This design enables BiMediX2 to generate precise and context-aware responses in either English or Arabic, depending on the user prompt, ensuring flexibility and inclusivity.
Key to BiMediX2’s performance is its modular and efficient training approach. LoRA adapters (Hu et al. (2021)) are utilized to fine-tune the language model while maintaining computational efficiency and minimizing resource demands. The projector is simultaneously fine-tuned to optimize image-text alignment in a medical context. Furthermore, the system is supported by a robust data generation framework, where a comprehensive English data corpus is translated into Arabic using GPT-4o. A random subset of this translation is meticulously verified by bilingual medical experts to ensure clinical relevance and linguistic accuracy. This pipeline enables BiMediX2 to excel in a wide range of medical tasks, including report generation, radiology analysis, pathology insights, and ophthalmological assessments, all within a unified, bilingual, and multimodal framework.
BiMed-V Bilingual and Multimodal Dataset Creation
The BiMed-V dataset is a comprehensive bilingual and multimodal instruction set comprising of 1.6M samples, developed to enhance medical image-text alignment and multimodal understanding. It incorporates a diverse range of publicly available datasets, such as PMC-OA (Lin et al. (2023)), Rad-VQA (Lau et al. (2018)), Path-VQA (He et al. (2020)) and SLAKE (Liu et al. (2021)) complemented by custom-curated data. We also curated 163k VQA samples by repurposing the LLaVA-Med (Li et al. (2023)) 60K-IM dataset, which were tailored to align with real-world medical queries. Furthermore, over 10k samples from the LLaVA-Med pretraining dataset were reformatted into interactive conversations using the Llama 3.1 70B model. A subset of the PMC-OA dataset with short question-answer pairs and multiple-choice questions were added to enhance the dataset’s diversity. Training splits of Rad-VQA, Path-VQA and SLAKE, which typically feature concise answers, were restructured into more detailed responses using the same Llama 3.1 70B model, enhancing the dataset’s depth and usability for complex tasks.
A unique feature of BiMed-V is its bilingual support, facilitated by a multimodal instruction set comprising 326k samples across various medical imaging modalities. This includes 163k Arabic-language samples generated via a comprehensive translation framework (see Fig. 2). English datasets were translated into Arabic using GPT-4o, with rigorous verification of a random subset by bilingual medical experts to ensure clinical relevance and linguistic precision. This hybrid approach balances automation and expert validation, significantly reducing reliance on human medical domain experts while maintaining data quality. Additionally, the inclusion of text-based clinical data from BiMediXv1 (Pieri et al. (2024)) ensures the dataset retains robust language understanding capabilities while expanding its multimodal medical proficiency. This extensive dataset forms the foundation for advanced medical image-text alignment and conversational multimodal applications.
Bilingual and Multimodal Medical Instruction Tuning
To enable BiMediX2’s robust capabilities in both bilingual and multimodal medical tasks, we employ a two-stage training process that ensures precise alignment of visual and textual representations while adapting the language model for complex medical instruction tasks.
Stage 1 Training: Medical Concept Alignment
In the first stage, we finetune the Projector alone to align visual embeddings to the language embedding space. The training utilizes a comprehensive dataset of 467k image-caption pairs sourced from the LLaVA-Med (Li et al. (2023)) pretraining dataset. These pairs span diverse medical imaging modalities and captions that describe clinically relevant features.
Stage 2 Training: Multimodal Medical Instruction Alignment
The second stage finetunes the LoRA(Hu et al. (2021)) adapters within the language model, enhancing its ability to process and generate multimodal medical instructions. For this, we utilize our BiMed-V 1.6M bilingual multimodal instruction set, which comprises meticulously crafted English and Arabic prompts paired with corresponding visual and textual responses. This dataset enables the model to learn nuanced instructions across a wide array of medical domains, from radiology to pathology, in a bilingual context.
Through these two stages, BiMediX2 achieves seamless integration of bilingual and multimodal capabilities, enabling it to deliver accurate and context-aware medical insights in both English and Arabic, tailored to a variety of clinical scenarios.
3 Experiments
In the literature, evaluating medical language models predominantly involves multiple-choice question-answering tasks, with accuracy as the performance metric. We employed the EleutherAI (Gao et al. (2024)) evaluation framework for evaluating text-only medical tasks following (Pieri et al. (2024)). For our bilingual and multimodal evaluations, we utilized several benchmarks and tools. We leveraged our BiMed-MBench for assessing the correctness, helpfulness, relevance and accuracy of the model in a bilingual medical multimodal context. Additionally, we evaluated the model on VQA using datasets such as Path-VQA, SLAKE, and Rad-VQA. For report generation, we used the MIMIC-CXR (Johnson et al. (2019)) dataset, and for report summarization, we used the MIMIC-III (Johnson et al. (2016)) dataset. These evaluations were conducted using the MultiMedEval (Royer et al. (2024)) toolkit, ensuring a comprehensive assessment of the model’s capabilities in both textual and visual medical tasks.
Medical LLM Benchmarks
@@ 3 Experiments
PubMedQA (Jin et al. (2019)) is a question-answering dataset derived from biomedical research papers on PubMed. Given a question sourced from the title of a research paper and a context from the abstract, the task is to provide an answer in the form of ’yes’, ’no’, or ’maybe’, mirroring the conclusion of the paper. Building on prior research, our examination focuses on the PQA-L subset of PubMedQA, consisting of 500 manually annotated QA pairs sourced from PubMed. Answering these questions with accuracy requires thorough reasoning over the biomedical contexts and quantitative data analysis.
MedMCQA (Pal et al. (2022)) is a multiple-choice dataset constructed from questions featured in the Indian AIIMS and NEET PG medical entrance exams. It covers a broad spectrum of medical subjects, evaluating both professional domain knowledge and language comprehension. The dataset includes a test set of 4,183 questions, each with four options.
MedQA (Jin et al. (2021)) is a dataset comprising multiple-choice questions from medical board examinations in the US, Mainland China, and Taiwan. It features two types of questions: concise sentences to assess specific knowledge and extended paragraphs that detail a patient’s condition. Our analysis concentrates on the English portion (USMLE), sourced directly from the National Board of Medical Examiners (NBME), responsible for the assessments required for medical licensure in the U.S., containing 1,273 samples for testing purposes. Engaging with this benchmark involves multiple stages of reasoning and evidence retrieval.
USMLE (Han et al. (2023)) USMLE Self-assessment Test, derived from the United States Medical Licensing Examination (USMLE) self-assessment for Step 1, Step 2, and Step 3, is composed of multiple-choice questions. We employ the version provided by the authors of MedAlpaca, which specifically omits image-based questions and comprises 325 test samples.
The Medical MMLU (Hendrycks et al. (2020)) (Massive Multitask Language Understanding) is a group of six datasets that compiles 1,089 test questions with four options related to different medical subjects. The six MMLU datasets are Clinical Knowledge (Cli-KG), College Biology (C-Bio), College Medicine (C-Med), Medical Genetics (Med-Gen), Professional Medicine (Pro-Med) & Anatomy (Ana).
UPHILL OpenQA (Kaur et al. (2023)) The UPHILL dataset is designed to evaluate the factual accuracy of LLMs in handling health-related queries in English embedded with varying degrees of presuppositions. Our analysis specifically targeted the zero-shot models’ abilities to process and respond to false claims (factually incorrect or debunked by reliable sources) across a spectrum of presupposition levels, a crucial test given the potential real-world implications of misinformation and inaccuracies in the health domain. In this context, the accuracy refers to the model’s effectiveness in accurately refuting false health-related claims at different presupposition levels.
Medical VLM Benchmarks
BiMed-MBench: We introduced the first bilingual GPT-4o-based medical LMM benchmark, consisting of 286 medical queries spanning various medical imaging modalities. The test dataset includes conversational interactions and detailed descriptions for modalities such as chest X-rays (CXR), MRI, histology, gross and CT scans derived from LLaVA-Med (Li et al. (2023)). GPT-4o evaluates the correctness of model responses based on the provided image context and caption. The reference prediction, serving as the upper-bound answer for the teacher model, is generated using GPT-4 (Achiam et al. (2023)). For Arabic evaluations, the ground truth is derived from the upper-bound, translated by GPT-4o, and all test samples are comprehensively verified by medical experts for quality assurance. GPT-4o evaluates the responses by comparing the candidate LMM against the upper-bound for English queries and against the Arabic ground truth for Arabic queries. The evaluation assesses the helpfulness, relevance, accuracy, and level of detail of the responses while maintaining the user language. The scoring is conducted on a scale of 1 to 10, with higher scores indicating better performance. Additionally, GPT-4o provides detailed explanations of the evaluation to offer deeper insights into the models’ performance. Finally, we compute and compare the relative scores of the two candidates.
Visual Question Answering (VQA): To evaluate a model’s performance in Visual Question Answering, three datasets are used: Path-VQA, SLAKE, and Rad-VQA. These datasets include a mix of open-ended and close-ended questions, designed to test the model’s ability to interpret and respond to queries based on medical images. This evaluation ensures the model’s effectiveness in handling a variety of visual and textual inputs.
Report Generation: The MIMIC-CXR (Johnson et al. (2019)) dataset, which includes de-identified radiology reports associated with chest X-rays, is utilized to assess the model’s performance in generating the findings section of medical reports. For this task, the input consists of one or more radiology images related to a case, followed by a prompt such as, “<<<image>>> <<<image>>> Please caption this scan with findings and impressions.” This setup evaluates the model’s ability to generate accurate and coherent medical reports based on visual inputs.
Report Summarization: We conduct an evaluation of report summarization using the MIMIC-III (Johnson et al. (2016)) dataset. The task involves generating the impressions section of a radiology report based on the findings section. To achieve this, the model is provided with the free-text findings along with a task prompt, “Summarize the findings”. This assessment focuses on the model’s ability to distill comprehensive medical information into a concise and precise summary, which is essential for clear and effective medical communication.
4 Results
LLM Evaluation
The performance of BiMediX2 and other models on various medical benchmarks is presented in Fig 3 and Tab 2. Our BiMediX2 70B achieved the highest average score of 84.6%, outperforming other models such as GPT-4 (82.9%) and Llama-3-Med42-70B (83.0%). BiMediX2 70B exhibited superior average performance across multiple datasets, including Medical MMLU, MedMCQA, MedQA, USMLE, and PubMedQA, demonstrating its strong understanding of medical contexts.

### v2

@@ Abstract
We introduce BiMediX2, a bilingual (Arabic-English) Bio-Medical EXpert Large Multimodal Model that supports text-based and image-based medical interactions. It enables multi-turn conversation in Arabic and English and supports diverse medical imaging modalities, including radiology, CT, and histology. To train BiMediX2, we curate BiMed-V, an extensive Arabic-English bilingual healthcare dataset consisting of 1.6M samples of diverse medical interactions. This dataset supports a range of medical Large Language Model (LLM) and Large Multimodal Model (LMM) tasks, including multi-turn medical conversations, report generation, and visual question answering (VQA). We also introduce BiMed-MBench, the first Arabic-English medical LMM evaluation benchmark, verified by medical experts. BiMediX2 demonstrates excellent performance across multiple medical LLM and LMM benchmarks, achieving state-of-the-art results compared to other open-sourced models. On BiMed-MBench, BiMediX2 outperforms existing methods by over 9% in English and more than 20% in Arabic evaluations. Additionally, it surpasses GPT-4 by approximately 9% in UPHILL factual accuracy evaluations and excels in various medical VQA, report generation, and report summarization tasks. Our trained models, instruction set, and source code are available at – https://github.com/mbzuai-oryx/BiMediX2
BiMediX2: Bio-Medical EXpert LMM for Diverse Medical Modalities
Sahal Shaji Mullappilly1thanks: Equal Contribution,  Mohammed Irfan Kurpath [^1] footnotemark: 1,  Sara Pieri1  Saeed Yahya Alseiari5,  Shanavas Cholakkal6,  Khaled Aldahmani3,4,  Fahad Khan1,2  Rao Anwer1,  Salman Khan1, Timothy Baldwin1,  Hisham Cholakkal1 1Mohamed Bin Zayed University of Artificial Intelligence (MBZUAI), 2Linköping University 3Shaikh Tahnoon bin Mohammed Medical City (STMC), 4Tawam Hospital 5Sheikh Shakhbout Medical City (SSMC), 6Govt Medical College Kozhikode
@@ 1 Introduction
Figure 1: Performance comparison on BiMed-MBench. The comparison is conducted across different tasks and modalities, including CT, MRI, CXR, Histology, and Gross, along with their Arabic counterparts (CT_ar, MRI_ar, CXR_ar, Histology_ar, and Gross_ar). Each axis represents the performance score for a specific category, highlighting BiMediX2’s superior performance across diverse tasks and modalities in both English and Arabic.
@@
Meditron Chen et al. (2023)
Med42 Christophe et al. (2024)
OpenBioLLM Ankit Pal (2024)
Llama3.1 Meta (2024)
BiMediXv1 Pieri et al. (2024)
Model MTC RS RG Rad Oph Path Micro UM LLM+VLM Bil (Ar)
LLaVA-pp Rasheed et al. (2024)
MiniGPT-Med Alkhaldi et al. (2024)
MAIRA-2 Bannur et al. (2024)
BioMedGPT Zhang et al. (2024)
LLaVA-Med Li et al. (2023)
Dragonfly-Med Chen et al. (2024)
@@
Table 1: Comparison of tasks and modalities addressed by recent medical LLMs and VLMs. Abbreviations: MTC (Multi-turn conversation), RS (Report Summarization), RG (Report Generation), Rad (Radiology), Oph (Ophthalmology), Path (Pathology), Micro (Microscopic), UM (Unified Model: Single model checkpoint for all downstream tasks), LLM+VLM (Unified LLM + VLM), Bil (Ar) (Bilingual Arabic capabilities).
Recently, medical Large Language Models (LLMs) and medical Large Multimodal Models (LMMs) have shown promising results as conversational assistants for improving accessibility to quality medical advice. However, most medical Vision-Language Models (VLMs) referred to interchangeably as medical LMMs in this paper often compromise their text-based understanding (i.e., medical LLM performance) when integrating multimodal capabilities, making it challenging to interact seamlessly with users (see LLM+VLM column in Tab. 1). This is particularly challenging when users initially ask general medical queries in text format and later follow up with questions related to user-provided medical images introduced mid-conversation to provide additional supporting information.
In addition to supporting diverse LLM and LMM tasks, such as multi-turn conversations and report generation, it is also desirable to support various medical image modalities, such as radiology and pathology, in a unified model. However, as shown in Tab. 1, state-of-the-art medical LMMs, such as MiniGPT-Med Alkhaldi et al. (2024), are either restricted to a limited set of medical image modalities, such as radiology, or require separately fine-tuned models for each downstream task (e.g., BiomedGPT Zhang et al. (2024)), posing significant challenges for real-world deployment.
Moreover, advancements in medical LLMs and LMMs remain predominantly English-centric, leaving significant gaps for non-English-speaking populations, particularly in languages like Arabic, spoken by over 400 million people. Unlike most state-of-the-art medical LLMs, BiMediX Pieri et al. (2024) (referred to as BiMediXv1 in this paper for clarity) supports Arabic-English bilingual interactions and offers diverse text-based interaction capabilities, such as multi-turn conversations, report summarization, and question-answering. However, it is limited to text modality and lacks medical image understanding capabilities.
To address the aforementioned limitations of existing medical LMMs, we propose BiMediX2, a Bio-Medical EXpert Large Multimodal Model that supports diverse medical tasks and modalities while also facilitating seamless user interactions in both English and Arabic.
1.1 Contributions
Figure 2: BiMediX2: Overall Architecture Our model is designed for medical image analysis and bilingual multi-turn conversations. Medical images are processed through a Vision Encoder and aligned with a Projector, while the text inputs are tokenized using the default tokenizer. The resulting tokens are then passed into the language model (Meta Llama 3.1) to generate responses in the prompted language. We only train the language model using LoRA adapters, while the projector is finetuned for medical image-text alignment. BiMediX2 follows a two-stage training pipeline. Stage-1 aligns medical visual concepts using 467K image-caption pairs. Stage-2 performs multimodal medical instruction tuning with our proposed BiMed-V 1.6M bilingual instructions comprising both image-text and text-only medical instructions.
We introduce BiMediX2 an Arabic-English bilingual medical LMM that supports broader spectrum of text-based and multimodal medical tasks, including Multi-Turn Conversations (MTC), Report Summarization (RS), Report Generation (RG), Radiology (Rad), Ophthalmology (Oph), Pathology (Path), Microscopic Analysis (Micro), in a Unified Model (UM) with LLM and VLM capabilities (LLM+VLM) as shown in Tab. 1. It supports a wide range of imaging modalities, including Chest X-rays, CT scans, MRIs, Histology slides, and Gross pathology images (see Fig. 1). The key contributions of our work can be summarized as: (i) We curated a comprehensive Arabic-English bilingual healthcare specialized instruction set named BiMed-V comprising over 1.6M instructions including text-only and image-text instructions across diverse medical image modalities.
@@ 1.1 Contributions
(ii) We introduce the first Arabic-English bilingual medical LMM evaluation benchmark named BiMed-MBench, consisting of 386 medical queries in English and Arabic across various medical image modalities, fully verified by medical experts.
(iii) We instruction-tune Llama3.1 LLM with our BiMed-V to obtain the first Arabic-English bilingual medical LMM achieving state-of-the-art results among open-sourced models on various VLM evaluation benchmarks across diverse medical image modalities while also excelling on medical LLM evaluation benchmarks in both English and Arabic. This is achieved through our multistage training pipeline and carefully curated instruction set, which balances text-only and multimodal medical instructions.
(iv) We perform comprehensive evaluation of our model across 12 diverse evaluation benchmarks. In addition to achieving promising results on our bilingual BiMed-MBench, BiMediX2 achieves excellent results for visual question answering on Path-VQA, SLAKE and Rad-VQA evaluation benchmarks, and report generation task on MIMIC CXR and report summarization on MIMIC-III benchmarks. Additionally, BiMediX2 achieves favorable results on diverse medical LLM benchmarks, such Medical MMLU, MedMCQA, MedQA, USMLE, and PubMedQA datasets, and it also shows robustness in discerning and correcting misinformation in medical context on the UPHILL OpenQA Evaluation.
@@
Fig 1 illustrates the performance of SOTA medical LMMs on our BiMed-MBench evaluation. BiMediX2 achieves SOTA results on BiMed-MBench, with over a 9% improvement in English evaluations and more than a 20% improvement in Arabic evaluations. Furthermore, it excels in medical Visual Question Answering, Report Generation, and Report Summarization tasks. Similarly, BiMediX2 outperforms GPT-4 by more than 8% on the USMLE benchmark and by more than 9% in UPHILL factual accuracy evaluations.
@@ 2 BiMediX2
The architecture of BiMediX2 is designed to facilitate seamless integration of medical image analysis and bilingual multi-turn conversations. At its core, (see Fig 2) the model employs a Vision Encoder Radford et al. (2021) to process a diverse array of medical imaging modalities, including chest X-rays, CT scans, MRIs, histology slides, and gross pathology images. This visual data is aligned with textual inputs through a dedicated Projector, ensuring accurate and contextually rich medical image-text mapping following Liu et al. (2023). As shown in Fig 2 we use the ‘<<image>>’ token as a place holder to encode the visual features for multimodal medical instructions. For text based medical data the inputs are processed using a standard tokenizer, transforming them into the language embedding space of Llama 3.1 Meta (2024). This design enables BiMediX2 to generate precise and context-aware responses in either English or Arabic, depending on the user prompt and supports multimodal interactions while preserving the medical LLM capabilities.
Key to BiMediX2’s performance is its modular and efficient training approach. LoRA adapters Hu et al. (2021) are utilized to fine-tune the language model while maintaining computational efficiency and minimizing resource demands. The projector is simultaneously fine-tuned to optimize image-text alignment in a medical context. Furthermore, the system is supported by a robust data generation framework, where a comprehensive English data corpus is translated into Arabic using GPT-4o. A random subset of this translation is meticulously verified by bilingual medical experts to ensure clinical relevance and linguistic accuracy. This pipeline enables BiMediX2 to excel in medical tasks, including report generation, radiology analysis, pathology insights, and ophthalmological assessments, in a unified, bilingual, and multimodal framework.
2.1 BiMed-V: Multimodal, Bilingual Dataset
The BiMed-V dataset is a comprehensive bilingual and multimodal instruction set comprising of 1.6M samples, developed to enhance medical image-text alignment and multimodal understanding. It incorporates a diverse range of publicly available datasets, such as PMC-OA Lin et al. (2023), Rad-VQA Lau et al. (2018), Path-VQA He et al. (2020) and SLAKE Liu et al. (2021) complemented by custom-curated data. We also curated 163k VQA samples by repurposing the LLaVA-Med Li et al. (2023) 60K-IM dataset. Furthermore, over 10k samples from the LLaVA-Med pretraining dataset were reformatted into interactive conversations using the Llama 3.1 70B model. A subset of the PMC-OA dataset with short question-answer pairs and multiple-choice questions were added to enhance the dataset’s diversity. Training splits of Rad-VQA, Path-VQA and SLAKE, which typically feature concise answers, were restructured into more detailed responses using the same Llama 3.1 70B model, enhancing the dataset’s depth and usability for complex tasks.
A unique feature of BiMed-V is its bilingual support, facilitated by a multimodal open-ended instruction set comprising 326k samples across various medical imaging modalities. This includes 163k Arabic-language samples generated via a comprehensive translation framework (see Fig. 5). English datasets were translated into Arabic using GPT-4o, with verification of a random subset by bilingual medical experts to ensure clinical relevance and linguistic precision. This hybrid approach balances automation and expert validation, significantly reducing reliance on human medical domain experts while maintaining data quality (Please refer A.4 for more details on translation framework and expert validation). Additionally, the inclusion of text-based clinical data from BiMediXv1 Pieri et al. (2024) ensures the dataset retains robust language understanding capabilities while expanding its multimodal medical proficiency. This extensive dataset forms the foundation for advanced medical image-text alignment and conversational multimodal applications. For dataset composition, see Section A.7.